two sailors decide to have an unusual race.Each has a power boat that travels at 10 m/s in still water.However,the river,of width 100 m,flows at 5 m/s. sailor A will go to a point directly across the starting point and return. sailor B will travel 100 m downstream and then return .who wins?

Grade:11

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Dear student,
Time taken by sailor A = 100/10 + 100/10 = 20 seconds
Time taken by sailor B = 100/15 + 100/5 = 80/3 seconds

hence, Sailor A wins.
….….….….….….….….….….….….….….….….….….….….….….…...
Thanks.